The Jones Company has just completed the third year of a​ five-year MACRS recovery period for a piece of equipment it originally purchased for $298,000.

a. What is the book value of the​ equipment?

b. If Jones sells the equipment today for $175,000 and its tax rate is 35%​, what is the​ after-tax cash flow from selling​ it?

c. Just before it is about to sell the​ equipment, Jones receives a new order. It can take the new order if it keeps the old equipment. Is there a cost to taking the order and if​ so, what is​ it? Explain.​ (Assume the new order will consume the remainder of the​ machine's useful​ life.)

Note​:

Assume that the equipment is put into use in year 1.
Answer #1

a. What is the book value of the​ equipment?

The Jones Company has just completed the third year of a​ five-year MACRS recovery period for a piece of equipment it originally purchased for $298,000.

Total depreciation within 3 years of 5 years MACRS, D = 20% + 32% + 19.2% = 71.2%

Hence, book value of the equipment = Initial purchase price x (1 - D) = 298,000 x (1 - 71.2%) =  85,824.00

--------------------------------

b. If Jones sells the equipment today for $175,000 and its tax rate is 35%​, what is the​ after-tax cash flow from selling​ it?

Gain = Sale price - Book Value = 175,000 -  85,824.00 =  89,176

The​ after-tax cash flow from selling​ it = Sale price - tax on gain = Sale price - tax rate x Gain on sale= 175,000 - 35% x 89,176 =  143,788.40

-----------------------------------

c. Just before it is about to sell the​ equipment, Jones receives a new order. It can take the new order if it keeps the old equipment. Is there a cost to taking the order and if​ so, what is​ it? Explain.​ (Assume the new order will consume the remainder of the​ machine's useful​ life.)

Note​: Assume that the equipment is put into use in year 1.

Yes, there is an opportunity cost involved in taking the order. Jones lose the opportunity to earn an after tax cash flows of $ 143,788.40 due to disposal of the old asset. By taking the order, Jones would have to forego the opportunity to sell this old equipment and thus forego the opportunity to earn $ 143,788.40, immediately.